# EqDrive: Efficient Equivariant Motion Forecasting with Multi-Modality for Autonomous Driving

### v2

@@ V Experiments
Model
minADE at 3s
minFDE at 3s
Miss Rate
Number of Parameters
Training Time
LaneGCN[9]
0.71
1.08
0.1
3.7M
8hrs w/ 4xTitanX
DenseTNT[23]
0.75
1.05
0.1
1.1M
5hrs w/ 8x2080Ti
HiVT-128[10]
0.661
0.969
0.092
2.5M
56.5hrs w/ 1x3060Ti
@@ V-B Metrics for Evaluation
    where $\mathbf{y}_{t}^{h}$ represents the predicted position at time $t$ from the $h$-th trajectory, $\mathbf{y}_{t}$ indicates the ground truth position at time $t$, and $\tau$ is the prediction horizon. In our case, $\tau$ can be $10$ (for 1s), $20$ (for 2s), or $30$ (for 3s) given our 10Hz sampling frequency.
@@
    Here, $\tau$ can take values $10$, $20$, or $30$.
@@ V-C Implementation Details
Description
Symbol
Value
Length of input sequence
$T_{\text{in}}$
20
Length of output sequence
$T_{\text{out}}$
30
Number of total agents in the scene
$A$
4
Number of lane centerlines
$L$
10
Number of coordinates per centerline
$K$
100
Number of prediction heads
$H$
6
Number of repeats on the feature learning layers
$Q$
20
Size of all hidden dimension
$hidden\_dim$
64
Loss Component Weight
$\beta$
0.5

### v3

@@ V Experiments
Model minADE at 3s minFDE at 3s Miss Rate Number of Parameters Training Time
LaneGCN[9] 0.71 1.08 0.1 3.7M 8hrs w/ 4xTitanX
DenseTNT[23] 0.75 1.05 0.1 1.1M 5hrs w/ 8x2080Ti
HiVT-128[10] 0.661 0.969 0.092 2.5M 56.5hrs w/ 1x3060Ti
@@ V-B Metrics for Evaluation
    where $\mathbf{y}_{t}^{h}$ represents the predicted position at time $t$ from the $h$-th trajectory, $\mathbf{y}_{t}$ indicates the ground truth position at time $t$, and $\tau$ is the prediction horizon. In our case, $\tau$ can be 101010 (for 1s), 202020 (for 2s), or 303030 (for 3s) given our 10Hz sampling frequency.
@@
    Here, $\tau$ can take values 101010, 202020, or 303030.
@@ V-C Implementation Details
Description Symbol Value
Length of input sequence $T_{\text{in}}$ 20
Length of output sequence $T_{\text{out}}$ 30
Number of total agents in the scene $A$ 4
Number of lane centerlines $L$ 10
Number of coordinates per centerline $K$ 100
Number of prediction heads $H$ 6
Number of repeats on the feature learning layers $Q$ 20
Size of all hidden dimension $hidden\_dim$ 64
Loss Component Weight $\beta$ 0.5